# Multiverse: Transactional Memory with Dynamic Multiversioning

### v1

@@ Abstract.
conference: Arxiv; 2025; journalyear: 2025copyright: noneccs: Computing methodologies Concurrent algorithms
@@ 1. Introduction
Software transactional memory (STM) is a synchronization mechanism that allows users to execute sequences of memory accesses as atomic transactions. STMs make atomicity easy but not necessarily fast. STMs typically perform read-only transactions using optimistic synchronization. In such STMs (herlihy1993transactional; dice2006transactional; lev2009anatomy; ramalhete2024scaling; felber2008dynamic; fraser2007concurrent; memoryprinciples; ennals2006software; shavit1995software), memory addresses are typically associated with version numbers that are used to determine whether a transaction’s reads are consistent. If not, a transaction aborts and retries. For workloads where transactions are small (accessing few addresses) and contention is low, TMs can typically achieve good performance. On the other hand, STMs often struggle to handle transactions that read a large number of addresses that are frequently updated. This type of transaction is very likely to repeatedly abort.
@@
A classical solution in data structures and databases for supporting large read-only operations is to utilize multiversioned concurrency control (MVCC), which allows one to take atomic snapshots of memory (afek1993atomic; wu2017empirical). Much of the early work on MVCC focused on maintaining consistent versions and avoiding an unbounded number of versions per address. Many of these MVCC designs which focus only on providing atomic snapshots are typically expensive. Recent work has made significant practical improvements to MVCC (blelloch2024verlib; wei2021constant).
@@
There have been some attempts to combine multiversioning with STM (lu2013generic; perelman2011smv; riegel2006snapshot). These existing approaches typically guarantee the weaker correctness conditions of snapshot isolation (berenson1995critique) or serializability (bernstein1987concurrency), as opposed to (the stronger) opacity (tm-book). Without opacity, aborted transactions are allowed to observe inconsistent state which can lead to various problems (dice2014pitfalls). Furthermore, maintaining multiple versions can incur substantial overhead, reducing performance in the case where versioning is not required. In such cases, traditional unversioned TMs are preferable.
@@
In this work we present Multiverse, a new opaque STM that leverages recent advancements in MVCC (blelloch2024verlib) to enable read-only transactions that access a large number of addresses to commit even in the presence of many concurrent updates that would otherwise cause potentially unbounded aborts. Unlike prior multiversioned STMs, our primary design goal is to ensure that our unversioned transactions are competitive with the fastest unversioned STM (DCTL at present (ramalhete2024scaling)) while also ensuring that our versioned transactions are fast as possible subject to the constraint that they do not add substantial overhead to unversioned transactions. As Figure 1 shows, Multiverse can outperform the fastest unversioned STM even for workloads with very few range queries (RQs), where versioning is not always necessary.
@@ 2.2. Correctness
Snapshot isolation (SI) (berenson1995critique), serializability, strict serializability (bernstein1987concurrency) and the stronger opacity (tm-book) are common correctness conditions used in TM. A history of committed transactions is serializable if it is equivalent to some serial history. Strict serializability further requires that the history maintains the real-time order of transactions. Opacity requires that the history of all transactions (including aborted ones) be equivalent to some sequential history. In other words, opacity requires that all transactions must observe consistent state. As discussed in (dice2014pitfalls), one should care about opacity since it prevents various problems. In particular, without opacity, one loses even single-threaded invariants since transactions could observe inconsistent state and continue running. The weaker SI, intuitively, allows transactions to perform (consistent) reads in the past, but write in the present, which can be difficult to use correctly. Opacity is most common in TM.
@@ 3.3.1. Transitioning between TM Modes
3.4. Correctness
@@
In this section we provide a brief intuitive correctness argument for opacity. Our unversioned transactions ensure consistency via the versioned locks. On each TM access, the lock versions are validated against the transaction’s read clock. Any inconsistency immediately causes an abort. This validation ensures that a transaction only observes writes from other committed transactions; a similar argument applies to versioned transactions in Mode Q that version addresses. Our versioned transactions ensure consistency following an approach similar to Verlib. During a traversal of a version list, the timestamp of each individual version is validated against the transaction’s versioned timestamp. This ensures that for each address read, the transaction only observes versions that were written by other committed transactions. Furthermore, failing to find a suitable version immediately causes an abort. The use of timestamp based validation makes it easy to see that an equivalent sequential order would order transactions based on their read clocks / versioned timestamps. In the case that two transaction share the same read clock / versioned timestamp they can only both commit if they are disjoint, thus they can be ordered arbitrary relative to each other in an equivalent sequential order. Likewise, since transactions abort immediately upon encountering an inconsistency, even aborted transactions cannot observe inconsistent state.
@@ 4.5. Memory Management
In section 2.3 of the TL2 paper, the authors mention the problem of memory reclamation but they only discuss an issue where we want to free objects that exist within the write-set of a transaction. Their proposed reclamation scheme only works if the issue can be caught during read-set validation, which will never occur in our example. This problem is also briefly discussed in (blelloch2025tlf) and the authors similarly avoid it via EBR. Alternatively, we could require read-only transactions perform revaluation or we could utilize an allocator where the internal state is accessed via transactions. Both alternatives have significant performance drawbacks.
@@ 5. Evaluation
We implemented Multiverse in C++. The implementation is publicly available [^1] The code can be found at: https://gitlab.com/Coccimiglio/mvcc_tm. The code was compiled with GCC 10.3.0 with an optimization level of -O2. We compare against several existing STMs which also guarantee opacity. Specifically TL2 (dice2006transactional), DCTL (ramalhete2024scaling), NOrec (dalessandro2010norec) and TinySTM (felber2008dynamic). These STMs are described in Section 6. For TL2 and TinySTM we use the author’s public implementations. For NOrec we used same implementation as in (brown2019cost). While DCTL has no public implementation, the authors have a public implementation of a similar algorithm (ramalhete2021efficient) which we could easily adapt into DCTL. We plan to make our code public after publication. We used the same benchmark as (brown2020non) for our evaluation. We ran all experiments on a single AMD EPYC 7662 processor which has 64 cores and 128 hardware threads. All results report the average of 5 trials. The measurement period of each trial is 20 seconds. In this section we focus on a representative sample of our results using an (a,b)-tree. In Appendix A we show other data structures including an internal AVL tree, an external binary search tree and a hashmap.
@@
For TL2 we use the GV4 global clock implementation. DCTL requires specifying the number of aborts before falling back to a starvation free mode for which we use 100 (the maximum used in (ramalhete2024scaling)). For Multiverse we use $\mathcal{K}_{1}$=100, $\mathcal{K}_{2}$=16, $\mathcal{K}_{3}$=28, $\mathcal{S}$=10, $\mathcal{L}$=10 and $\mathcal{P}$=10%. For both Multiverse and DCTL we use the same linear backoff as in (ramalhete2021efficient). Unmentioned parameters use default values.
@@ 1 Updater
There are some application benchmarks for TM such as STAMP (minh2008stamp), TPC-C (tpcc) and YCSB (cooper2010benchmarking) however these benchmarks have various problems which make them not interesting if we care about performance for workloads with long running queries in the presence of frequent updates. STAMP was designed to demonstrate the advantages of first generation TM which had no practical support for large queries. TPC-C is almost an append-only workload despite defining many transaction types, more than 90% of transactions are simple, append-only new-order or payment transactions. In popular research implementations of TPC-C, like DBx1000’s, the default index type is a hash table, which can not reasonably be used to implement RQs, and ad-hoc solutions are hard-coded instead (yu2014staring). YCSB-A is a rather boring nearly read-only workload. There are no row insertions or deletes. One can interpret our benchmark as a generalization of YCSB-A in which we not only update rows, but also insert and delete rows. Issues with TPC-C and YCSB are discussed in detail in section 7 of (avni2016persistent).
@@ 0 Updaters
Power Consumption is another interesting metric to consider. We compare the energy efficiency of the TMs via perf by measuring the power consumption of the CPU package in joules (specifically measuring the energy-pkg hardware event). Note that it is not possible to measure the energy consumption of a specific program or core (khan2018rapl; raffin2024dissecting). Figure 10 shows average throughput per joule of energy for the same (a,b)-tree from row two of Figure 6. Multiverse is able to leverage increased memory usage to efficiently support RQs resulting in up to 50x improved energy efficiency compared to the next best TM.
@@ 6. Related Work
Transactional Locking II (TL2) (dice2006transactional) is one of the most well known STMs. TL2 is an opaque unversioned STM that relies on a global clock and per-address versioned locks. By default, TL2 uses buffered writes and deferred commit time locking. Transactions that write increment the global clock at commit time after locking the write set. Numerous optimizations of TL2 exist, especially new approaches for managing the global clock. The most recent of these optimizations was presented in a new STM, Deferred Clock Transactional Locking (DCTL), (ramalhete2024scaling). DCTL uses encounter time locking and increments the global clock only on aborts. It has a starvation free mode where a single transaction can become irrevocable. Multiple non-irrevocable transactions can execute concurrently with a single irrevocable transaction. TLRW (dice2010tlrw) is an STM that relies on byte-locks which are a form of read-write locks. TLRW is unversioned but it does support irrevocable transactions. It has been shown to be competitive with TL2 for single chip machines in some workloads. Unfortunately TLRW has no public implementation. TinySTM (felber2008dynamic) is another well known opaque unversioned STM. Like TL2, it relies on a global clock and per-address versioned locks. TinySTM uses encounter time locking. No Ownership Records (NORec) (dalessandro2010norec) is an unversioned STM that does not use per-address versioned locks. It uses a single global sequence lock, commit time locking and value based validation. These STMs do not have proper support for long running read-only transactions.
@@
Verlib (blelloch2024verlib) is the state of the art MVCC mechanism. Notably, Verlib can sometimes avoid adding indirection when updating addresses. Verlib was incorporated into a sort of restricted STM that can be used only for implementing optimistic data structures in (blelloch2025tlf). Unlike Multiverse  this restricted TM does not guarantee opacity. Other non-opaque multiversion STMs have also been proposed (riegel2006snapshot; lu2013generic). Selective Multi-Versioning (SMV) (perelman2011smv) is an opaque STM that uses an approach somewhat similar to Verlib. We would have liked to compare against SMV (perelman2011smv), but it is implemented in Java and heavily relies on garbage collection. To our knowledge there is no public non-Java implementation and it is not straightforward to implement it in C++, since we would need to solve the memory management problem. SMV uses commit time locking. In SMV updates always add new versions but old versions can be quickly removed. Active transactions in SMV append a descriptor to a global list. A descriptor has a timestamp and keeps references to the old versions of objects that the transaction modified to prevent their reclamation by the garbage collector. Unlike Verlib which uses a numeric global timestamp, SMV uses the list of descriptors to serve as a global clock which requires different synchronization.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: ; ; $c$s: Computing methodologies Concurrent algorithms
@@ 1. Introduction
Software transactional memory (STM) is a synchronization mechanism that allows users to execute sequences of memory accesses as atomic transactions. STMs make atomicity easy but not necessarily fast. STMs typically perform read-only transactions using optimistic synchronization. In such STMs (Herlihy and Moss, 1993; Dice et al., 2006; Lev et al., 2009; Ramalhete and Correia, 2024; Felber et al., 2008; Fraser and Harris, 2007; Guerraoui and Kapałka, 2010; Ennals and Ennals, 2006; Shavit and Touitou, 1995), memory addresses are typically associated with version numbers that are used to determine whether a transaction’s reads are consistent. If not, a transaction aborts and retries.
For workloads where transactions are small (accessing few addresses) and contention is low, TMs can typically achieve good performance. On the other hand, STMs often struggle to handle transactions that read a large number of addresses that are frequently updated. This type of transaction is very likely to repeatedly abort.
@@
A classical solution in data structures and databases for supporting large read-only operations is to utilize multiversioned concurrency control (MVCC), which allows one to take atomic snapshots of memory (Afek et al., 1993; Wu et al., 2017). Much of the early work on MVCC focused on maintaining consistent versions and avoiding an unbounded number of versions per address. Many of these MVCC designs which focus only on providing atomic snapshots are typically expensive. Recent work has made significant practical improvements to MVCC (Blelloch and Wei, 2024; Wei et al., 2021).
@@
There have been some attempts to combine multiversioning with STM (Lu and Scott, 2013; Perelman et al., 2011; Riegel et al., 2006). These existing approaches typically guarantee the weaker correctness conditions of snapshot isolation (Berenson et al., 1995) or serializability (Bernstein et al., 1987), as opposed to (the stronger) opacity (Harris et al., 2022). Without opacity, aborted transactions are allowed to observe inconsistent state which can lead to various problems (Dice et al., 2014). Furthermore, maintaining multiple versions can incur substantial overhead, reducing performance in the case where versioning is not required. In such cases, traditional unversioned TMs are preferable.
@@
In this work we present Multiverse, a new opaque STM that leverages recent advancements in MVCC (Blelloch and Wei, 2024) to enable read-only transactions that access a large number of addresses to commit even in the presence of many concurrent updates that would otherwise cause potentially unbounded aborts. Unlike prior multiversioned STMs, our primary design goal is to ensure that our unversioned transactions are competitive with the fastest unversioned STM, Deferred Clock Transactional Locking (DCTL) at present (Ramalhete and Correia, 2024), while also ensuring that our versioned transactions are fast as possible subject to the constraint that they do not add substantial overhead to unversioned transactions. As Figure 1 shows, Multiverse can outperform the fastest unversioned STM even for workloads with very few range queries (RQs), where versioning is not always necessary.
@@ 2.2. Correctness
Snapshot isolation (SI) (Berenson et al., 1995), serializability, strict serializability (Bernstein et al., 1987) and the stronger opacity (Harris et al., 2022) are common correctness conditions used in TM. A history of committed transactions is serializable if it is equivalent to some serial history. Strict serializability further requires that the history maintains the real-time order of transactions. Opacity requires that the history of all transactions (including aborted ones) be equivalent to some sequential history. In other words, opacity requires that all transactions must observe consistent state. As discussed in (Dice et al., 2014), one should care about opacity since it prevents various problems. In particular, without opacity, one loses even single-threaded invariants since transactions could observe inconsistent state and continue running. The weaker SI, intuitively, allows transactions to perform (consistent) reads in the past, but write in the present, which can be difficult to use correctly. Opacity is most common in TM.
@@ 3.3.1. Transitioning between TM Modes
3.4. Correctness and Progress
Theorem 3.1.
Multiverse guarantees weak progressiveness and opacity.
Weak Progressiveness
In Multiverse, a conflict will occur if a transaction attempts to validate a lock that is already locked. A conflict will also occur if a transaction attempts to validate a lock with a version that is greater than or equal to the transaction’s read clock. Both scenarios can only arise as a result of a concurrent (update) transaction. When two transactions conflict Multiverse does not guarantee that one of the transactions will commit, meaning it is not strongly progressive. However, since aborts only occur due to conflicts, Multiverse does guarantee weak progressiveness.
Opacity
@@ 3.4. Correctness and Progress
In Multiverse it is relatively straightforward to see that a history in which the TM mode is fixed is opaque. Unversioned transactions in Multiverse ensure consistency via the versioned locks. On each TM access, the lock versions are validated against the transaction’s read clock. Any inconsistency immediately causes an abort. This validation ensures that a transaction only observes writes from other committed transactions; a similar argument applies to versioned transactions in Mode Q that version addresses. Versioned transactions in Multiverse ensure consistency following an approach similar to Verlib. During a traversal of a version list, the timestamp of each individual version is validated against the transaction’s versioned timestamp. This ensures that for each address read, the transaction only observes versions that were written by other committed transactions. Furthermore, failing to find a suitable version immediately causes an abort. The use of timestamp based validation makes it easy to see that an equivalent sequential order would order transactions based on their read clocks / versioned timestamps. In the case that two transaction share the same read clock / versioned timestamp they can only both commit if they are disjoint, thus they can be ordered arbitrary relative to each other in an equivalent sequential order. Likewise, since transactions abort immediately upon encountering an inconsistency, even aborted transactions cannot observe inconsistent state.
A history in which the TM mode changes is slightly more complicated due to the fact that the local mode of a transaction can differ from the global mode. Before discussing this case, it is worth noting that since the core synchronization of update transactions in Multiverse is a TL2 style synchronization, the classic case based TL2 opacity proof (Brown and Ravi, 2017) can be translated to apply to Multiverse which would completely remove the notion of the Multiverse modes (or any similar construct).
Dealing with the TM modes becomes significantly easier when considering histories that are partitioned based on the global TM mode. Recall the following facts regarding mode transitions: All mode transitions are atomic. The transition from Mode Q to Mode QtoU can be performed by worker threads, however, since the global mode is a monotonically increasing integer, only one thread will succeed the CAS to advance the mode in this case. All other mode transitions are centralized to the background thread and occur sequentially in a fixed order. Thus, the global mode will always progress in the cyclic order of Q, QtoU, U, and UtoQ.
Furthermore, the (integer) value of the local mode of any thread is at most one less than the global mode. If some thread $t$ had a local mode that was more than one away from the local mode then when the background thread iterated over the thread local data it must have found the local mode of $t$ was one behind the current mode and still went through with incrementing the global mode. However, in this scenario the background thread would reiterate over the thread local data meaning the local mode of $t$ would at worst remain one behind the global mode. [^1] The background thread iterating over the announcement array is sufficient to ensure that this invariant holds if we assume that threads never leave the system and rejoin. If threads are allowed to leave and rejoin then additional synchronization would be needed. Based on the above, it is easy to show that there does not exist a history containing a Mode U reader when the global mode is Mode Q or Mode QtoU.
With these invariants in mind, the opacity guarantees of Multiverse can be proven via induction on the number of global TM mode changes. I provide an intuitive summary of this proof below for a history $H$ containing $j>=0$ mode changes. It is clear that a prefix of $H$ containing 0 mode changes is opaque since this will be a history containing only transactions in Mode Q which is effectively equivalent to DCTL. Now assume that a prefix of $H$ containing up to $k<n$ mode changes is opaque. The induction requires showing that a prefix of $H$ containing $k+1$ mode changes is opaque. This can be split into 4 cases based on the global mode following the $k$-th mode change: (1): After the $k$-th mode change the TM is in Mode Q meaning after the $(k+1)$-th mode change the global TM is in Mode QtoU. (2): After the $k$-th mode change the TM is in Mode QtoU meaning after the $(k+1)$-th mode change the global TM is in Mode U. (3): After the $k$-th mode change the TM is in Mode U meaning after the $(k+1)$-th mode change the global TM is in Mode UtoQ. (4): After the $k$-th mode change the TM is in Mode UtoQ meaning after the $(k+1)$-th mode change the global TM is in Mode Q. Note that this is simply one case per mode transition.
For each case consider what changes between the modes. For case (1), after transitioning to Mode QtoU all new or retrying writers version (meaning the writers operate as if they are in Mode U) but readers remain in Mode Q. For case (2), after transitioning to Mode U, the history can contain Mode U readers. Writers version in all modes except Mode Q and while the TM is in Mode U, the local mode of any thread cannot be Mode Q. For case (3), after transitioning to Mode UtoQ all new or retrying readers will no longer rely on the invariant of writers versioning (meaning the readers operate as if they are in Mode Q) but all active and new writers will still version. Finally, for case (4), after transitioning to Mode Q writers will no longer version but there will be no readers with local Mode U.
Each of the above 4 cases can itself be broken down into several sub-cases considering all possible pairs of conflicting transactions with different local modes. The opacity proof for each sub-case would follow the same approach as in Claim 5 of Appendix A of (Brown and Ravi, 2017) with the additional possibility of establishing a contradiction based on the difference between the local mode of a transaction and the global mode.
@@ 4.5. Memory Management
In section 2.3 of the TL2 paper, the authors mention the problem of memory reclamation but they only discuss an issue where we want to free objects that exist within the write-set of a transaction. Their proposed reclamation scheme only works if the issue can be caught during read-set validation, which will never occur in our example. This problem is also briefly discussed in (Blelloch et al., 2025) and the authors similarly avoid it via EBR. Alternatively, we could require read-only transactions perform revaluation or we could utilize an allocator where the internal state is accessed via transactions. Both alternatives have significant performance drawbacks.
@@ 5. Evaluation
We implemented Multiverse in C++. The implementation is publicly available (Coccimiglio et al., 2025) [^2] The code can be found at: https://zenodo.org/records/18099743. The code was compiled with GCC 10.3.0 with an optimization level of -O2. We compare against several existing STMs which also guarantee opacity. Specifically TL2 (Dice et al., 2006), DCTL (Ramalhete and Correia, 2024), NOrec (Dalessandro et al., 2010) and TinySTM (Felber et al., 2008). These STMs are described in Section 6. For TL2 and TinySTM we use the author’s public implementations. For NOrec we used same implementation as in (Brown and Ravi, 2017). While DCTL has no public implementation, the authors have a public implementation of a similar algorithm (Ramalhete et al., 2021) which we could easily adapt into DCTL. We used the same benchmark as (Brown et al., 2020) for our evaluation. We ran all experiments on a single AMD EPYC 7662 processor which has 64 cores and 128 hardware threads. All results report the average of 5 trials. The measurement period of each trial is 20 seconds. In this section we focus on a representative sample of our results using an (a,b)-tree. In Appendix A we show other data structures including an internal AVL tree, an external binary search tree and a hashmap.
@@
For TL2 we use the GV4 global clock implementation. DCTL requires specifying the number of aborts before falling back to a starvation free mode for which we use 100 (the maximum used in (Ramalhete and Correia, 2024)). For Multiverse we use $\mathcal{K}_{1}$=100, $\mathcal{K}_{2}$=16, $\mathcal{K}_{3}$=28, $\mathcal{S}$=10, $\mathcal{L}$=10 and $\mathcal{P}$=10%. For both Multiverse and DCTL we use the same linear backoff as in (Ramalhete et al., 2021). Unmentioned parameters use default values.
@@ 1 Updater
There are some application benchmarks for TM such as STAMP (Minh et al., 2008), TPC-C (Transaction Processing Performance Council, 2010) and YCSB (Cooper et al., 2010) however these benchmarks have various problems which make them not interesting if we care about performance for workloads with long running queries in the presence of frequent updates. STAMP was designed to demonstrate the advantages of first generation TM which had no practical support for large queries. TPC-C is almost an append-only workload despite defining many transaction types, more than 90% of transactions are simple, append-only new-order or payment transactions. In popular research implementations of TPC-C, like DBx1000’s, the default index type is a hash table, which can not reasonably be used to implement RQs, and ad-hoc solutions are hard-coded instead (Yu et al., 2014). YCSB-A is a rather boring nearly read-only workload. There are no row insertions or deletes. One can interpret our benchmark as a generalization of YCSB-A in which we not only update rows, but also insert and delete rows. Issues with TPC-C and YCSB are discussed in detail in section 7 of (Avni and Brown, 2016).
@@ 0 Updaters
Power Consumption is another interesting metric to consider. We compare the energy efficiency of the TMs via perf by measuring the power consumption of the CPU package in joules (specifically measuring the energy-pkg hardware event). Note that it is not possible to measure the energy consumption of a specific program or core (Khan et al., 2018; Raffin and Trystram, 2025). Figure 10 shows average throughput per joule of energy for the same (a,b)-tree from row two of Figure 6. Multiverse is able to leverage increased memory usage to efficiently support RQs resulting in up to 50x improved energy efficiency compared to the next best TM.
@@ 6. Related Work
Transactional Locking II (TL2) (Dice et al., 2006) is one of the most well known STMs. TL2 is an opaque unversioned STM that relies on a global clock and per-address versioned locks. By default, TL2 uses buffered writes and deferred commit time locking. Transactions that write increment the global clock at commit time after locking the write set. Numerous optimizations of TL2 exist, especially new approaches for managing the global clock. The most recent of these optimizations was presented in a new STM, Deferred Clock Transactional Locking (DCTL), (Ramalhete and Correia, 2024). DCTL uses encounter time locking and increments the global clock only on aborts. It has a starvation free mode where a single transaction can become irrevocable. Multiple non-irrevocable transactions can execute concurrently with a single irrevocable transaction. TLRW (Dice and Shavit, 2010) is an STM that relies on byte-locks which are a form of read-write locks. TLRW is unversioned but it does support irrevocable transactions. It has been shown to be competitive with TL2 for single chip machines in some workloads. Unfortunately TLRW has no public implementation. TinySTM (Felber et al., 2008) is another well known opaque unversioned STM. Like TL2, it relies on a global clock and per-address versioned locks. TinySTM uses encounter time locking. No Ownership Records (NORec) (Dalessandro et al., 2010) is an unversioned STM that does not use per-address versioned locks. It uses a single global sequence lock, commit time locking and value based validation. These STMs do not have proper support for long running read-only transactions.
@@
Verlib (Blelloch and Wei, 2024) is the state of the art MVCC mechanism. Notably, Verlib can sometimes avoid adding indirection when updating addresses. Verlib was incorporated into a sort of restricted STM that can be used only for implementing optimistic data structures in (Blelloch et al., 2025). Unlike Multiverse this restricted TM does not guarantee opacity. Other non-opaque multiversion STMs have also been proposed (Riegel et al., 2006; Lu and Scott, 2013). Selective Multi-Versioning (SMV) (Perelman et al., 2011) is an opaque STM that uses an approach somewhat similar to Verlib. We would have liked to compare against SMV (Perelman et al., 2011), but it is implemented in Java and heavily relies on garbage collection. To our knowledge there is no public non-Java implementation and it is not straightforward to implement it in C++, since we would need to solve the memory management problem. SMV uses commit time locking. In SMV updates always add new versions but old versions can be quickly removed. Active transactions in SMV append a descriptor to a global list. A descriptor has a timestamp and keeps references to the old versions of objects that the transaction modified to prevent their reclamation by the garbage collector. Unlike Verlib which uses a numeric global timestamp, SMV uses the list of descriptors to serve as a global clock which requires different synchronization.